OpenStudy (anonymous):

How do I factor 4n^2-8n+3???Please help?

jimthompson5910 (jim_thompson5910):

the first coefficient is 4 the last term is 3 multiply them to get 4*3 = 12 ------------------------------------- Now list out the possible ways to multiply to 12 12 = 1*12 12 = 2*6 12 = 3*4 12 = -1*(-12) 12 = -2*(-6) 12 = -3*(-4) which of these pairs of factors add to -8 ?

OpenStudy (anonymous):

-2 and 6...But I think I keep writing 8n factored wrong,I keep putting 4n^2-2-6+3...Is that part correct??

jimthompson5910 (jim_thompson5910):

4n^2-8n+3 should break into 4n^2-2n - 6n+3

jimthompson5910 (jim_thompson5910):

because -2 and -6 a) multiply to 12 AND b) add to -8

jimthompson5910 (jim_thompson5910):

now you factor 4n^2-2n - 6n+3 by grouping

jimthompson5910 (jim_thompson5910):

4n^2-2n - 6n+3 (4n^2-2n) + (-6n+3) 2n(2n - 1) - 3(2n - 1) What's next?

OpenStudy (anonymous):

I egt n(4n-2)+2(3n+1),how do I get to the right answer for this part.I keep getting the same thing but its wrong.

jimthompson5910 (jim_thompson5910):

do you see how I got 2n(2n - 1) - 3(2n - 1) ?

OpenStudy (anonymous):

Not really,Could you please explain it?

jimthompson5910 (jim_thompson5910):

you see how I got 4n^2-2n - 6n+3 right?

OpenStudy (anonymous):

yeah,that part I understand.

jimthompson5910 (jim_thompson5910):

ok in the next step, I grouped the terms into pairs like this (4n^2-2n) + (-6n+3)

jimthompson5910 (jim_thompson5910):

basically the first two terms get in one set of parenthesis the next two get the next set of parenthesis

jimthompson5910 (jim_thompson5910):

from there, I factored the GCF of the first group out of the first group the GCF of 4n^2 - 2n is 2n So I factored 2n from 4n^2 - 2n to get 4n^2 - 2n = 2n(2n - 1)

jimthompson5910 (jim_thompson5910):

Then I factored -3 from -6n + 3 to get -6n+3 = -3(2n - 1)

jimthompson5910 (jim_thompson5910):

so that's how I went from (4n^2-2n) + (-6n+3) to 2n(2n - 1) - 3(2n - 1)

OpenStudy (anonymous):

Ok,so the next part should look like (2n-3)(2n-1).And that would be the answer?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

Thank you very much :)

jimthompson5910 (jim_thompson5910):

you can expand that out and combine like terms you should get the original expression back again that's one way to check the answer

jimthompson5910 (jim_thompson5910):

yw